# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ E.2 Relative convexity and smoothness between functionals
$\int\left\langle v(x),\int\Big{(}\big{(}\nabla^{2}W(x-y)-\beta\nabla^{2}V(x)\big{)}v(x)-\nabla^{2}W(x-y)v(y)\Big{)}\mathrm{d}\mu_{k}(y)\right\rangle\mathrm{d}\mu_{k}(x)\leq 0.$
For example, choosing $W(x)=\frac{1}{2}\|x\|_{2}^{2}$, then $\nabla^{2}W=I_{d}$ and $\mathcal{F}$ is $\beta$-smooth relative to $\mathcal{G}$ as long as $\nabla^{2}V\succeq\frac{1}{\beta}I_{d}$.
@@ Appendix F Bregman proximal gradient scheme
The first step of our analysis is to show that this scheme is equivalent with
@@
Let $\mu_{0}\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, $\tau\leq\frac{1}{\beta}$ and $\mathcal{F}(\mu)=\mathcal{G}(\mu)+\mathcal{H}(\mu)$ with ~$\tilde{\mathcal{H}}_{\mu_{k}}$ convex on $L^{2}(\mu)$ and $\mathcal{G}$ $\beta$-smooth relative to $\phi$ along $t\mapsto\big{(}(1-t)\mathrm{Id}+t\tilde{\mathrm{T}}_{k+1}\big{)}_{\#}\mu_{k}$. Then, for all $\mathrm{T}\in L^{2}(\mu_{k})$,
@@
$\mathcal{F}(\mu_{k+1})\leq\mathcal{F}(\mu_{k})-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{Id},\tilde{\mathrm{T}}_{k+1}).$ (106)
Additionally, let $\alpha\geq 0$, $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and suppose that $\phi_{\mu}$ satisfies 1. If $\mathcal{G}$ is $\alpha$-convex relative to $\phi$ along $t\mapsto\big{(}(1-t)\mathrm{Id}+t\mathrm{T}_{\phi_{\mu_{k}}}^{\mu_{k},\nu}\big{)}_{\#}\mu_{k}$, then for all $k\geq 1$,
@@
We verify now that Proposition 22 can be applied for mirror schemes of interest. Salim et al. [101, Lemma 2] showed that it holds for the Wasserstein proximal gradient of potentials, i.e. with $\phi(\mu)=\int\frac{1}{2}\|\cdot\|_{2}^{2}\ \mathrm{d}\mu$ and $\mathcal{G}(\mu)=\int U\ \mathrm{d}\mu$ with $U$ (strictly) convex. We extend their result for $\mathcal{G}(\mu)=\int U\ \mathrm{d}\mu$ and $\phi(\mu)=\int V\ \mathrm{d}\mu$ for $V$ strictly convex and $U$ $\beta$-smooth relative to $V$.
@@
Let $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, $\mathcal{F}(\mu)=\int U\ \mathrm{d}\mu$, $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ with $V$ strongly convex and $U$ $\beta$-smooth relative to $V$, and $\mathrm{T}=\nabla V^{*}\circ(\nabla V-\tau\nabla U)$. Assume $\tau<\frac{1}{\beta}$, then $\mathrm{T}_{\#}\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$.
Proof.
The proof of the lemma is inspired from [101, Lemma 2]. The goal is to apply [4, Lemma 5.5.3], which requires to show that $\mathrm{T}$ is injective almost everywhere and that $|\det\nabla\mathrm{T}|>0$ almost everywhere. See Section H.11 for the proof. ∎
To apply Proposition 23, we also need $\mathcal{H}$ to be convex. Let $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, and denote $\rho$ its density w.r.t the Lebesgue measure. For $\mathcal{H}(\mu)=\int f\big{(}\rho(x)\big{)}\ \mathrm{d}x$ where $f:\mathbb{R}\to\mathbb{R}$ is $C^{1}$ and satisfies $f(0)=0$, $\lim_{x\to 0}xf^{\prime}(x)=0$ and $x\mapsto f(x^{-d})x^{d}$ is convex and non-increasing on $\mathbb{R}_{+}$, then by [109, Theorem 4.2], $\mathcal{H}$ is convex along curves $\mu_{t}=\big{(}(1-t)\mathrm{S}+t\mathrm{T})_{\#}\mu$ obtained with $\mathrm{S}$ and $\mathrm{T}$ with positive definite Jacobians. This is the case e.g. for $f(x)=x\log x$, for which $\mathcal{H}$ corresponds to the negative entropy.
@@
By Remark 3, to be able to apply the three-point inequality to have the descent lemma, we need $\mathcal{H}$ to be convex along $\big{(}(1-t)\tilde{\mathrm{T}}_{k+1}+t\mathrm{Id}\big{)}_{\#}\mu_{k}$ and along $\big{(}(1-t)\tilde{\mathrm{T}}_{k+1}+t\mathrm{T}_{\phi_{\mu_{k}}}^{\mu_{k},\nu}\big{)}_{\#}\mu_{k}$ for the convergence. Assuming the covariances matrices are full rank, $\tilde{\mathrm{T}}_{k+1}$ is affine and its gradient is invertible. Moreover, by Proposition 13, $\mathrm{T}_{\phi_{\mu_{k}}}^{\mu_{k},\nu}=\nabla u\circ\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})$ for $\nabla u$ an OT map between $\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})_{\#}\mu_{k}$ and $\nu$. Since everyone is Gaussian, and $\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})(x)=\Lambda^{-1}x$ is affine, is has a positive definite Jacobian. Thus, using [109, Theorem 4.2], we can conclude that we can apply Proposition 23.
Closed-form for Gaussian.
@@
Since $\nabla V(x)=\Lambda^{-1}x$, and $\mu_{k}=\mathcal{N}(m_{k},\Sigma_{k})$, we obtain for all $x\in\mathbb{R}^{d}$,
$\mathrm{S}_{k+1}(x)=\Lambda\big{(}\Lambda^{-1}x-\tau\Sigma^{-1}(x-m)\big{)}=x-\tau\Lambda\Sigma^{-1}(x-m).$ (109)
@@
Moreover, if $\Lambda$ and $\Sigma_{k+1}$ commute, this is equivalent with
@@ Appendix G Additional details on experiments
G.1 Solving the general scheme
@@
Except for the case $\phi(\mu)=\int V\ \mathrm{d}\mu$ where $\nabla_{\mathrm{W}_{2}}\phi(\mu)=\nabla V$ does not depend on $\mu$, one cannot in general invert $\nabla_{\mathrm{W}_{2}}\phi(\mu_{k+1})$ directly. It might be the case even for $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ if $\nabla V$ does not have an analytical inverse.
A practical workaround is to solve an implicit problem, see e.g. [121]. Here, we use the Newton-Raphson algorithm. Basically, suppose we have $\mu=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$. Then, the scheme is equivalent with
@@
$G_{j}(x_{1},\dots,x_{n})=\nabla_{\mathrm{W}_{2}}\phi\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}\right)(x_{j})-\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})(x_{j})+\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})(x_{j}).$ (118)
Write $\mathcal{G}(x_{1},\dots,x_{n})=\big{(}G_{1}(x_{1},\dots,x_{n}),\dots,G_{n}(x_{1},\dots,x_{n})\big{)}$. Then, we perform the following Newton iterations at each step:
$(x_{1}^{k+1},\dots,x_{n}^{k+1})=(x_{1}^{k},\dots,x_{n}^{k})-\gamma\big{(}J_{\mathcal{G}}(x_{1}^{k},\dots,x_{n}^{k})\big{)}^{-1}\mathcal{G}(x_{1}^{k},\dots,x_{n}^{k}).$ (119)
@@ G.2 Mirror descent of interaction energies
We detail in this Section the first experiment of Section 5. We aim at minimizing the interaction energy $\mathcal{W}(\mu)=\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$ for $W(z)=\frac{1}{4}\|z\|_{2}^{4}-\frac{1}{2}\|z\|_{2}^{2}$. It is well-known that the stationary solution of its gradient flow is a Dirac ring [26]. Since the stationary solution is translation invariant, we project the measures to be centered.
@@
with $p_{2}(t)=3t^{2}+1$. Thus, by [81, Remark 2], $W$ is $\beta$-smooth relative to $K_{4}(z)=\frac{1}{4}\|z\|_{2}^{4}+\frac{1}{2}\|z\|_{2}^{2}$ with $\beta=4$. Thus, using Proposition 21, $\tilde{\mathcal{W}}_{\mu}$ is $\beta$-smooth relative to $\phi_{\mu}(\mathrm{T})=\iint K\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ for all $\mu$, and we can apply Proposition 2.
@@
We also study the minimization of an interaction energy with an ill-conditioned kernel $W(z)=\frac{1}{4}(z^{T}\Sigma^{-1}z)^{2}-\frac{1}{2}z^{T}\Sigma^{-1}z$ where $\Sigma\in S_{d}^{++}(\mathbb{R})$ but is possibly badly conditioned, i.e. the ratio between the largest and smallest eigenvalues is big. In this case, the stationary solution becomes an ellipsoid instead of a ring. In our experiments, we take $\Sigma=\mathrm{diag}(100,0.1)$. For each scheme, we use $\mu_{0}=\mathcal{N}(0,0.25^{2}I_{2})$, $n=100$ particles and a step size of $\tau=0.1$.
@@
Similarly as in the not preconditioned case, using that $\nabla^{2}W(z)=2\Sigma^{-1}zz^{T}\Sigma^{-1}+(z^{T}\Sigma^{-1}z-1)\Sigma^{-1}$, we can show that
@@
For the sake of comparison, we also report on Figure 4 the trajectories of particles for the use of $K_{4}$ and $K_{4}^{\Sigma}$, as well as of the usual Wasserstein gradient descent and the preconditioned Wasserstein gradient descent obtained with $h^{*}(x)=\frac{1}{2}x^{T}\Sigma x$ (which is equivalent with the Mirror Descent with $\phi_{\mu}^{V}$ as Bregman potential and $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$). We observe almost the same trajectories as $K_{2}$, which would indicate that the target is also smooth compared to $\phi_{\mu}^{V}$.

### v2

@@ E.2 Relative convexity and smoothness between functionals
$\int\Big{\langle}v(x),\int\Big{(}\big{(}\nabla^{2}W\big{(}\mathrm{T}_{s}(x)-\mathrm{T}_{s}(y)\big{)}-\beta\nabla^{2}V\big{(}\mathrm{T}_{s}(x)\big{)}\big{)}v(x)$
$\quad-\nabla^{2}W\big{(}\mathrm{T}_{s}(x)-\mathrm{T}_{s}(y)\big{)}v(y)\Big{)}\mathrm{d}\ \mu_{k}(y)\Big{\rangle}\mathrm{d}\mu_{k}(x)\leq 0.$
For example, choosing $W(x)=\frac{1}{2}\|x\|_{2}^{2}$, then $\nabla^{2}W=I_{d}$ and $\mathcal{F}$ is $\beta$-smooth relative to $\mathcal{G}$ as long as $\nabla^{2}V\circ\mathrm{T}_{s}\succeq\frac{1}{\beta}I_{d}$ for any $s\in[0,1]$.
@@ Appendix F Bregman proximal gradient scheme
The first step of our analysis is to show that this scheme is equivalent to
@@
Let $\mu_{0}\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, $\tau\leq\frac{1}{\beta}$ and $\mathcal{F}(\mu)=\mathcal{G}(\mu)+\mathcal{H}(\mu)$. Consider the iterates of the Bregman proximal gradient scheme (104), equivalently (105). Let $k\geq 0$. Assume ~$\tilde{\mathcal{H}}_{\mu_{k}}$ is convex on $L^{2}(\mu_{k})$ and $\mathcal{G}$ $\beta$-smooth relative to $\phi$ along $t\mapsto\big{(}(1-t)\mathrm{Id}+t\tilde{\mathrm{T}}_{k+1}\big{)}_{\#}\mu_{k}$. Then, for all $\mathrm{T}\in L^{2}(\mu_{k})$,
@@
$\mathcal{F}(\mu_{k+1})\leq\mathcal{F}(\mu_{k})-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{Id},\tilde{\mathrm{T}}_{k+1}),$ (107)
i.e., the scheme decreases the objective at each iteration. Additionally, let $\alpha\geq 0$, $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and suppose that $\phi_{\mu}$ satisfies 1. If $\mathcal{G}$ is $\alpha$-convex relative to $\phi$ along $t\mapsto\big{(}(1-t)\mathrm{Id}+t\mathrm{T}_{\phi_{\mu_{k}}}^{\mu_{k},\nu}\big{)}_{\#}\mu_{k}$, then for all $k\geq 1$,
@@
We verify now that Proposition 24 can be applied for mirror schemes of interest. Salim et al. [109, Lemma 2] showed that it holds for the Wasserstein proximal gradient when using some potential energies, more precisely with $\phi(\mu)=\int\frac{1}{2}\|\cdot\|_{2}^{2}\ \mathrm{d}\mu$ and $\mathcal{G}(\mu)=\int U\ \mathrm{d}\mu$ with $U$ (strictly) convex. We extend their result for $\mathcal{G}(\mu)=\int U\ \mathrm{d}\mu$ and $\phi(\mu)=\int V\ \mathrm{d}\mu$ for $V$ strictly convex and $U$ $\beta$-smooth relative to $V$.
@@
Let $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, $\mathcal{G}(\mu)=\int U\ \mathrm{d}\mu$, $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ with $V$ strongly convex and $U$ $\beta$-smooth relative to $V$, and $\mathrm{T}=\nabla V^{*}\circ(\nabla V-\tau\nabla U)$. Assume $\tau<\frac{1}{\beta}$, then $\mathrm{T}_{\#}\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$.
Sketch of the proof.
The proof of the lemma is inspired from [109, Lemma 2]. We apply [5, Lemma 5.5.3], which requires to show that $\mathrm{T}$ is injective almost everywhere and that $|\det\nabla\mathrm{T}|>0$ almost everywhere. See Section H.13 for the full proof. ∎
To apply Proposition 25, we need $\mathcal{H}$ to be convex along some curve. We discuss here the convexity of the negative entropy along acceleration free curves. Let $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$, and denote $\rho$ its density w.r.t the Lebesgue measure. For $\mathcal{H}(\mu)=\int f\big{(}\rho(x)\big{)}\ \mathrm{d}x$ where $f:\mathbb{R}\to\mathbb{R}$ is $C^{1}$ and satisfies $f(0)=0$, $\lim_{x\to 0}xf^{\prime}(x)=0$ and $x\mapsto f(x^{-d})x^{d}$ is convex and non-increasing on $\mathbb{R}_{+}$, then by [117, Theorem 4.2], $\mathcal{H}$ is convex along curves $\mu_{t}=\big{(}(1-t)\mathrm{S}+t\mathrm{T})_{\#}\mu$ obtained with $\mathrm{S}$ and $\mathrm{T}$ with positive definite Jacobians. This is the case e.g. for $f(x)=x\log x$, for which $\mathcal{H}$ corresponds to the negative entropy.
By Remark 3, to be able to apply the three-point inequality (that is necessary to obtain the descent lemma), we actually only need $\mathcal{H}$ to be convex along $\big{(}(1-t)\tilde{\mathrm{T}}_{k+1}+t\mathrm{Id}\big{)}_{\#}\mu_{k}$ and along $\big{(}(1-t)\tilde{\mathrm{T}}_{k+1}+t\mathrm{T}_{\phi_{\mu_{k}}}^{\mu_{k},\nu}\big{)}_{\#}\mu_{k}$ for the convergence.
Gaussian target.
@@
Assuming the covariances matrices are full rank, $\tilde{\mathrm{T}}_{k+1}$ is affine and its gradient is invertible. Moreover, by Proposition 15, $\mathrm{T}_{\phi_{\mu_{k}}}^{\mu_{k},\nu}=\nabla u\circ\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})$ for $\nabla u$ an OT map between $\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})_{\#}\mu_{k}$ and $\nu$. Since each distribution is Gaussian, and $\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})(x)=\Lambda^{-1}x$ is affine, it has a positive definite Jacobian. Thus, using [117, Theorem 4.2], we can conclude that we can apply Proposition 25.
Closed-form for Gaussians.
@@
Since $\nabla V(x)=\Lambda^{-1}x$, and $\mu_{k}=\mathcal{N}(m_{k},\Sigma_{k})$, we obtain for $x_{k}\sim\mu_{k}$,
$\mathrm{S}_{k+1}(x_{k})=\Lambda\big{(}\Lambda^{-1}x_{k}-\tau\Sigma^{-1}(x_{k}-m)\big{)}=x_{k}-\tau\Lambda\Sigma^{-1}(x_{k}-m).$ (110)
@@
Moreover, if $\Lambda$ and $\Sigma_{k+1}$ commute, this is equivalent to
@@ Appendix G Additional details on experiments
G.1 Implementing the schemes
In this subsection, we sum up how to implement the different schemes in practice, given a finite number of particles. In all cases, we first sample $x_{1}^{(0)},\dots,x_{n}^{(0)}\sim\mu_{0}$, then we apply the scheme to $\hat{\mu}_{n}^{(k)}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}^{(k)}}$.
Mirror descent.
@@ G.1 Implementing the schemes
If $\phi(\mu)=\int V\ \mathrm{d}\mu$ with $\nabla V$ having an analytical inverse, the scheme can be implemented as
$\forall k\geq 0,\ \forall i\in\{1,\dots,n\},\ x_{i}^{(k+1)}=\nabla V^{*}\big{(}\nabla V(x_{i}^{(k)})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\hat{\mu}_{n}^{(k)})(x_{i}^{(k)})\big{)}.$ (118)
Except for this case, one cannot in general invert $\nabla_{\mathrm{W}_{2}}\phi(\mu_{k+1})\circ\mathrm{T}_{k+1}$ directly. A practical workaround is to solve an implicit problem, see e.g. [133]. Here, we use the Newton-Raphson algorithm. Suppose we have $\mu_{k}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}^{(k)}}$ and we are looking for $\mu_{k+1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$. Then, the scheme is equivalent to
@@
$G_{j}(x_{1},\dots,x_{n})=\nabla_{\mathrm{W}_{2}}\phi\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}\right)(x_{j})-\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})(x_{j}^{(k)})+\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})(x_{j}^{(k)}).$ (120)
Write $\mathcal{G}(x_{1},\dots,x_{n})=\big{(}G_{1}(x_{1},\dots,x_{n}),\dots,G_{n}(x_{1},\dots,x_{n})\big{)}$. Then, at each step $k$, we perform the following Newton iterations, starting from $(x_{1}^{(k)},\dots,x_{n}^{(k)})$:
$(x_{1}^{(k_{\ell+1})},\dots,x_{n}^{(k_{\ell+1})})=(x_{1}^{(k_{\ell})},\dots,x_{n}^{(k_{\ell})})-\gamma\big{(}J_{\mathcal{G}}(x_{1}^{(k_{\ell})},\dots,x_{n}^{(k_{\ell})})\big{)}^{-1}\mathcal{G}(x_{1}^{(k_{\ell})},\dots,x_{n}^{(k_{\ell})}).$ (121)
@@
Preconditioned gradient descent.
Plugging the empirical measure in (11), the preconditioned scheme can be implemented as
$\forall k\geq 0,\forall i\in\{1,\dots,n\},\ x_{i}^{(k+1)}=x_{i}^{(k)}-\tau\nabla h^{*}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\hat{\mu}_{n}^{(k)})(x_{i}^{(k)})\big{)}.$ (122)
@@ G.2 Mirror descent of interaction energies
We detail in this Section the first experiment of Section 5. We aim at minimizing the interaction energy $\mathcal{W}(\mu)=\frac{1}{2}\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$ for $W(z)=\frac{1}{4}\|z\|_{2}^{4}-\frac{1}{2}\|z\|_{2}^{2}$. It is well-known that the stationary solution of its gradient flow is a Dirac ring [27]. Since the stationary solution is translation invariant, we project the measures to be centered.
@@
with $p_{2}(t)=3t^{2}+1$. Thus, by [88, Remark 2], $W$ is $\beta$-smooth relative to $K_{4}(z)=\frac{1}{4}\|z\|_{2}^{4}+\frac{1}{2}\|z\|_{2}^{2}$ with $\beta=4$. Thus, using Proposition 23, $\tilde{\mathcal{W}}_{\mu}$ is $\beta$-smooth relative to $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\iint K\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ for all $\mu$, and we can apply Proposition 3.
@@
We also study the minimization of an interaction energy with an ill-conditioned kernel $W(z)=\frac{1}{4}(z^{T}\Sigma^{-1}z)^{2}-\frac{1}{2}z^{T}\Sigma^{-1}z$ where $\Sigma\in S_{d}^{++}(\mathbb{R})$ but is possibly badly conditioned, i.e. the ratio between the largest and smallest eigenvalues is large. In this case, the stationary solution becomes an ellipsoid instead of a ring. In our experiments, we take $\Sigma=\mathrm{diag}(100,0.1)$. For each scheme, we use $\mu_{0}=\mathcal{N}(0,0.25^{2}I_{2})$, $n=100$ particles and a step size of $\tau=0.1$.
@@
Similarly as in the case without preconditioning, using that $\nabla^{2}W(z)=2\Sigma^{-1}zz^{T}\Sigma^{-1}+(z^{T}\Sigma^{-1}z-1)\Sigma^{-1}$, we can show that
@@
For the sake of comparison, we also report on Figure 6 the trajectories of particles for the use of $K_{4}$ and $K_{4}^{\Sigma}$, as well as of the usual Wasserstein gradient descent and the preconditioned Wasserstein gradient descent obtained with $h^{*}(x)=\frac{1}{2}x^{T}\Sigma x$ (which is equivalent to the Mirror Descent with $\phi_{\mu}^{V}$ as Bregman potential and $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$). We observe almost the same trajectories as $K_{2}$, which would indicate that the target is also smooth compared to $\phi_{\mu}^{V}$.